# Inflationary Dynamics and Perturbations in Fractal Cosmology

### v1

@@ Abstract
We study inflationary dynamics within the framework of fractal cosmology, where spacetime exhibits a non-integer effective dimension, sourced through a relaxation of the cosmological principle. Using Friedmann and continuity equations, modified by an effective fractal dimension $D$; we derive generalized slow-roll parameters and examine their evolution for cubic, Starobinsky and Natural inflationary potentials. We then formulate a fractal extension of the Mukhanov-Sasaki equation by introducing an effective momentum term $k_{\text{eff}}$, arising from the fractal decomposition of the spatial Laplacian, that captures the geometric influence of fractal cosmology on scalar perturbations. This leads to corrections in the power spectrum and a scalar spectral index $n_{s}$ that depends explicitly on both the fractal dimension $D$ and a fractional scale $L$, which controls the strength of the fractal deformation. Comparison with the Planck 2018 data ($n_{s}=0.9649\pm 0.0042$) constrains the allowed range of $D$ ($2.7\lesssim D\lesssim 3$) depending on the cosmological and inflationary model assumed.
@@ 1 Introduction
Despite its successes, the assumption of large-scale homogeneity has come under scrutiny. Observations of galaxy clustering, filaments, and cosmic voids [5, 73, 120, 124, 102, 74, 67] suggest that the universe exhibits highly inhomogeneous, hierarchical structures across a broad range of scales. These features are naturally reminiscent of fractal geometry, raising the possibility that fractal cosmology [62, 27, 8, 104, 10, 89, 84, 91, 7, 106, 92, 29, 90, 76, 44, 13, 121, 22, 105] could offer an alternative or complementary framework to describe cosmic structure formation. In this approach, the geometry of spacetime and the associated cosmological dynamics acquire scale-dependent modifications [37], potentially leaving observable imprints on inflationary dynamics and perturbations.
Early discussions on a possible fractal geometry of the Universe and the observational debate about whether the galaxy distribution shows scale-invariant clustering; date back to the late 1980s and early 1990s. Seminal investigations include Aryal and Vilenkin’s analysis of fractal dimensions generated by inflationary scenarios [7] and the comprehensive reviews of the observational and theoretical status by Coleman [32], Pietronero [102] and Ribeiro [106]. Works like [104, 105, 37, 85] have studied quantum cosmology in the fractal setup (see for e.g. Chapter 7 of Jalalzedh & Moniz’s book [65]) Furthermore, there have been approaches where fractal inflationary models emerge from QG frameworks either through modes which have asymptotically safe gravity via running dimensions [76, 23, 44] or through Causal Dynamical Triangulations (CDT) [6, 91]. Moreover, fractal inflation can also emerge from non local QG theories [27, 26, 28]. These approaches provide concrete realizations of an effective dimension $D$ which motivates the phenomenological framework adopted in this work (see for e.g, the work by Khorrami, Mansouri et.al [69] and Mansouri & Nasseri [85].
@@
In this work, we investigate the dynamics of inflation in a fractal cosmological background  [^2] Although our framework introduces a fractional length scale $L$ and modifies the effective momentum structure (see Sec.4), this work should be clearly distinguished from fractional cosmology [53, 58, 115, 64, 77, 36, 118, 66, 117] based on fractional calculus or non-local fractional operators [98, 110, 55]. In particular, we do not employ non-local fractional derivatives (as compared to say works which use Riesz’s fractional derivative [64, 18, 97]) in the action [45, 47, 66] and model cosmology with a Orstein-Uhlenbeck like fractional differential equation [46]. Our approach is purely phenomenological and geometric, encoding derivatives from standard FLRW cosmology through an effective dimension $D$ and horizon thermodynamics. . By deriving modified slow-roll parameters, analyzing their evolution under different potentials (Cubic, Starobinsky ($R+R^{2}$) [93, 119, 38, 129, 83] and Natural [51, 50, 39]), and extending the perturbation theory to include fractal modifications, we aim to assess the viability of fractal inflationary models. In particular, we explore how the fractal dimension influences the duration of inflation, the behavior of perturbations, and the resulting predictions for the scalar spectral index [43, 101, 123, 31, 112, 88, 122]. Our results allow comparison with observational data, providing a means to constrain the fractal dimension and to evaluate the physical consistency of inflation within a fractal cosmology. Throughout our manuscript, we have treated the fractal dimension $D$ as a constant. This contrasts with models such as that of Abdalla, Afshori et. al [1] in which, $D$ evolves dynamically from near homogeneity ($D\simeq 3$) to a smaller asymptotic value, yielding a self-similar structure only in the late universe.
@@
2 A Review Fractal Cosmology
Before studying inflation in a fractal setup, it is necessary to establish the cosmological background in such a geometry. In fractal cosmology[62], the large-scale structure of spacetime is described by an effective dimension $D$, introducing corrections to the standard definitions of horizon area, entropy, and expansion dynamics[62]. In what follows, we review these modifications and derive the corresponding forms of the Friedmann and continuity equations that will form the basis of our inflationary analysis.
In standard cosmology, spacetime is treated as a smooth, differentiable manifold of integer dimension[17, 96, 128, 42, 78, 72]. However, in a fractal framework, this assumption is relaxed, and the effective dimensionality of spacetime becomes scale-dependent, reflecting the possibility that the microscopic structure of the universe may not be strictly three-dimensional. Various approaches to quantum gravity and spacetime discreteness, such as causal dynamical triangulations[6, 91] and asymptotic safety[76, 23, 44], suggest a dynamical reduction of the spectral dimension at small scales. Fractal cosmology provides a phenomenological framework for capturing such effects by modifying geometric quantities such as area, volume, and horizon entropy via an effective dimension $D\not=3$. These corrections, though small at large scales, can significantly alter the Friedmann equations and, in turn, the inflationary dynamics.
To begin with, recall that the line element of the FLRW universe [17, 96, 128, 42, 78, 72] is given by
@@ 2 A Review Fractal Cosmology
There are several nonequivalent ways to implement a fractal modifications of cosmology. The work of G. Calagani [27] proposed that phenomenological fractal measures and Lebesgue–Stieltjes integration which leads to a modified scaling of area/volume. There have been works where fractal-like spectral-dimension flows emerge in discrete quantum gravity approaches such as causal dynamical triangulations and asymptotic safety [76, 23, 44, 6, 91]. Furthermore, there have been thermodynamic derivations where the horizon entropy/area relations are deformed and the Friedmann equations follow from the first law on the apparent horizon (see for eg.[62]). In this work, we adopt the thermodynamic/cosmological horizon route as our working model.
@@
where $L$ denotes a characteristic fractional length scale associated with the underlying fractal microstructure. Throughout this work, we treat $L$ as a fixed microscopic scale of order Planck length. However in order to explain galaxy clustering, filaments and cosmic voids [5, 73, 120, 124, 102, 74, 67], you need a macroscopic value of $L$. The factor $(4\pi)^{(D-3)/3}$ ensures that the standard FLRW relations are recovered in the limit $D\rightarrow 3$.
@@ 3 Fractal Inflation Dynamics
Here $V(\phi)$ is the inflaton potential, and this inequality generalizes the standard slow-roll condition to a fractal spacetime.
@@
Substituting eqns.(16) and (17) into the continuity equation(14), we obtain the modified Klein-Gordon equation for the inflaton field:
@@
The presence of $D$ in the friction term indicates that the rate of damping of the inflaton’s motion depends directly on the effective fractal dimension; an increase in $D$ enhances the Hubble friction, thus delaying the end of inflation.
@@
$\ddot{H}=-\frac{4\pi GD^{2}}{9}\left(\frac{2GL^{2}\rho}{3}\right)^{\frac{D}{3}-1}\left[\left(\frac{D}{3}-1\right)\frac{(-DH\dot{\phi}^{2})}{\frac{\dot{\phi}^{2}}{2}+V(\phi)}-\dot{\phi}(DH\dot{\phi}+V_{,\phi})\right].$ (23)
@@
A sustained slow-roll inflation requires both $|\epsilon_{1}|<<1$ and $|\epsilon_{2}|<<1$ (note that $|\epsilon_{2}|<<1$ is a condition for slow-roll inflation, not inflation in general). The fractal background of the universe modifies $|\epsilon_{2}|$ to,
$\epsilon_{2}=\frac{D\dot{\phi}^{2}}{3\left[\frac{\dot{\phi}^{2}}{2}+V(\phi)\right]}[4-D]-2D-\frac{2V_{,\phi}}{H\dot{\phi}}=2\epsilon_{1}[4-D]-2D-\frac{2V_{,\phi}}{H\dot{\phi}}.$ (25)
@@
Figure 1: The above figure contains two plots. The left plot shows the evolution of $\epsilon_{1}$ with the number of e-folds, while the right plot shows the evolution of $\epsilon_{2}$ with the number of e-folds. Here we have chosen a linear potential $V=V_{0}(1+\frac{\phi}{10})$ and plotted for 4 cases; $D=1,2,3,4$. Here we have chosen the initial value of $\phi=10^{-2}$ and $\dot{\phi}=0$; corresponding to the scales expected for a constant offset monomial potential.
Figure 2: The above figure contains two plots therein. The left plot shows the evolution of $\epsilon_{1}$ with the number of e-folds, while the right plot shows the evolution of $\epsilon_{2}$ with the number of e-folds. Here we have chosen a cubic potential $V=V_{0}(1+\frac{\phi^{3}}{10})$ and plotted for 4 cases; $D=1,2,3,4$. Here we have chosen the initial value of $\phi=10^{-2}$ and $\dot{\phi}=0$, corresponding to the scales expected for a constant-offset monomial potential.
Figs.1 and 2 show the evolution of the slow roll parameter $\epsilon_{1}$ (given by the left plot) and $\epsilon_{2}$ (given by the right plot) with the number of e-folds in the case of a linear and cubic potential respectively. There are two key takeaways from the left hand sided plot of figs.1 and 2. One takeaway is that as $D$ increases, it takes more number of e-folds to end inflation, in other words, $N(\epsilon_{1})|_{D+1}-N(\epsilon_{1})|_{D}>0$; the second is that $N(\epsilon_{1})|_{D+1}-N(\epsilon_{1})|_{D}$ decreases as $D$ increases. Furthermore, from the right hand sided plot of figs.1 and 2, we infer that $\epsilon_{2}$ very rapidly tends to 0 as slow roll inflation starts. However, although the slow roll parameters show a dependence of the effective number of dimensions $D$, the role of $D$ in suppressing slow roll parameters subject to a monomial inflationary potential is relatively smaller than in Starobinsky and Natural inflation (discussed in later part of this manuscript). This leads us towards the direction that monomial potentials, showing less sensitivity to $D$, are more “preferred” in the fractal case than in standard $\Lambda$ CDM models, matching the recent observational results of the latest CMB and BAO data [9]. Moreover, the shape of the plots given by fig.1 and 2 don’t show much differences, at least with the initial conditions we have chosen, apart from scaling and exact numerical values.
Having analyzed monomial potentials such as cubic and linear, we now turn our attention to the Starobinsky ($R+R^{2})$ model of inflation [93, 119, 38, 129, 83], which arises from a purely geometric modification of gravity through the inclusion of an $R^{2}$ term in the action. Since this model is both theoretically well-motivated and tightly constrained by current CMB data, it provides an ideal benchmark for testing the robustness of the fractal framework against high-precision cosmological observations.
@@
$S=\frac{M_{Pl}^{2}}{2}\left(\frac{2GL^{2}\rho}{3}\right)^{1-\frac{D}{3}}\int d^{D+1}x\sqrt{-g}\left(R+\frac{R^{2}}{6M^{2}}\right).$ (28)
@@
$V(\phi)=\frac{3}{4}M^{2}M_{Pl}^{2}\left(\frac{2GL^{2}\rho}{3}\right)^{1-\frac{D}{3}}\left(1-e^{-\sqrt{\frac{D=1}{D}}\frac{\phi}{M_{Pl}}\left(\frac{2GL^{2}\rho}{3}\right)^{\frac{D}{6}-\frac{1}{2}}}\right)^{2}.$ (29)
@@
$\epsilon_{1}\simeq\frac{2(D-1)}{D}\frac{x^{2}}{(1-x)^{2}}\left[1-\frac{DH\dot{\phi}\phi}{\rho}\left(\frac{D}{3}-1\right)\right]^{2}.$ (30)
@@
$N\simeq\frac{D}{2(D-1)x}.$ (31)
@@
$\epsilon_{1}=\frac{D}{2(D-1)N^{2}}\left[1-\frac{DH\dot{\phi}\phi}{\rho}\left(\frac{D}{3}-1\right)\right]^{2},$ (32)
@@
$\epsilon_{2}=-\frac{4x}{(1-x)^{2}}\frac{D-1}{D}\left[1-\frac{DH\dot{\phi}\phi}{\rho}\left(\frac{D}{3}-1\right)\right]^{2}=-\frac{2}{N}\left[1-\frac{DH\dot{\phi}\phi}{\rho}\left(\frac{D}{3}-1\right)\right]^{2}.$ (33)
@@
In the fractal framework, the slow-roll parameter $\epsilon_{1}$ receives an additional multiplicative suppression $\left[1-\frac{DH\dot{\phi}\phi}{\rho}\left(\frac{D}{3}-1\right)\right]^{2}$ controlled by the effective dimension $D$, as shown in eqn.(32). For $D<3$ and $\phi<0$, this gravitational correction reduces $\epsilon_{1}$ and $|\epsilon_{2}|$ independently of the flattening of the inflaton potential. As a result, the phenomenological advantage of plateau models such as Starobinsky inflation becomes less pronounced: slow roll can be achieved without relying exclusively on extreme potential flatness. This leads to a partial degeneracy between potential-driven and geometry driven suppression mechanisms, rendering plateau models less uniquely preferred, though still fully consistent, within the fractal cosmology framework. This particular result goes in hand with the recent CMB and BAO data [9], providing a fractal perspective to the matter.
Note that, since the Higgs inflationary model [30, 108, 60, 109, 52] is also a plateau model like Starobinsky, we similarly expect that Higgs inflation is also less preferred in a fractal universe, once again going in hand with the recent CMB and BAO data [9]. A full analysis regarding the RG flow of $\lambda$ [116] and the subsequent reheating [113, 53] will be carried forward in future works.
We now turn to a qualitatively different class of inflationary scenarios, namely Natural Inflation [51, 50, 39]. In contrast to the Starobinsky potential, which originates from higher curvature corrections to gravity, Natural Inflation is driven by a pseudo Nambu Goldstone boson [57, 56, 54, 125, 126, 99, 130] whose flat potential is protected by an underlying shift symmetry.
@@
Let $f$ be the axion decay constant [54, 130] which serves to measure the extent of symmetry breaking and let $\Lambda$ set the energy scale for the symmetry breaking. The scalar potential for Natural Inflation is then given by a periodic form arising from the spontaneous breaking of a global symmetry, and is expressed as
@@
Hence, in the natural inflationary case, the only modification sourced by the fractal dimension $D$ is that of factor $\left(\frac{2GL^{2}\rho}{3}\right)^{1-\frac{D}{3}}$, which can effectively be thought of as changing the Planck mass as:-
$M_{Pl}\rightarrow M_{Pl}\left(\frac{2GL^{2}\rho}{3}\right)^{1-\frac{D}{3}}.$ (39)

### v2

# Inflation and Primordial Perturbations in Fractal Cosmology

@@ Abstract
We study inflationary dynamics within the framework of fractal cosmology, where space is characterized by an effective non-integer dimension $D$. In our work, fractal effects are sourced through thermodynamic modifications at the cosmological horizon. Using the modified Friedmann and continuity equations, we then derive the modified slow roll parameter and their evolution for linear, cubic, Starobinsky ($R+R^{2}$) and Natural inflationary potentials, showing that the slow roll parameters get suppressed for $D<3$. We further derive a fractal extension of the Mukhanov-Sasaki equation by introducing an effective momentum $k_{\text{eff}}$, which captures the modification of spatial Laplacian due to fractality. This leads to explicit corrections to the scalar power spectrum and the spectral index $n_{s}$, depending on both $D$ and a fractional length scale $L$. Confrontation with Planck 2018 data constrains the effective dimension to a best-fit range of $2.7\lesssim D\lesssim 3$ for the Starobinsky model. Furthermore, in the case of Natural Inflation, fractal corrections relax the usual requirement of super-Planckian axion decay constants, opening a phenomenologically viable parameter space inaccessible in the standard $3+1$ dimensional cosmology.
keywords:
Fractal cosmology , Inflationary dynamics , Perturbations , Horizon thermodynamics , Effective spacetime dimension
journal: Physics of the Dark Universe
[…]
[inst1]organization=Virtual Institute of Astroparticle Physics, addressline=75018 Paris, country=France
[…]
[inst2]organization=Departamento de Física, CMA-UBI, addressline=Universidade da Beira Interior, country=Portugal
[…]
[inst3]organization=Virtual Institute of Astroparticle Physics, addressline=75018 Paris, country=France
[…]
[inst4]organization=Department of Physics and Astronomy, Vanderbilt University, addressline=Nashville, TN 37235, country=USA
[…]
@@ 1 Introduction
Despite its successes, the assumption of large-scale homogeneity has come under scrutiny. Observations of galaxy clustering, filaments, and cosmic voids [4, 62, 100, 104, 84, 63, 60] suggest that the universe exhibits inhomogeneous, hierarchical structures across a broad range of scales. These features are naturally reminiscent of fractal geometry, raising the possibility that fractal cosmology [55, 21, 7, 86, 9, 74, 69, 76, 6, 89, 77, 25, 75, 64, 38, 12, 101, 16, 15, 88, 32, 52, 24] could offer an alternative or complementary framework to describe cosmic structure formation  [^1] Such an interpretation would implicitly require a macroscopic value of $L$. However, in the present work we restrict our analysis to the case where $L$ is treated as a microscopic parameter.. In this approach, the geometry of spacetime and the associated cosmological dynamics acquire scale-dependent modifications [34], potentially leaving observable imprints on inflationary dynamics and perturbations.
Early discussions on a possible fractal geometry of the Universe date back to the late 1980s and early 1990s. Seminal investigations include Aryal and Vilenkin’s analysis of fractal dimensions generated by inflationary scenarios [6] and the comprehensive reviews of the observational and theoretical status by Coleman [28], Pietronero [84] and Ribeiro [89]. Works like [86, 88, 34, 70] have studied quantum cosmology in the fractal setup (see for e.g. Chapter 7 of Jalalzedh & Moniz’s book [58]) Furthermore, there have been approaches where fractal inflationary models emerge from QG frameworks either through modes which have asymptotically safe gravity via running dimensions [64, 17, 38] or through Causal Dynamical Triangulations (CDT) [5, 76]. Moreover, fractal inflation can also emerge from non local QG theories [21, 20, 22]. These approaches provide concrete realizations of an effective dimension $D$ which motivates the phenomenological framework adopted in this work (see for e.g, the work by Khorrami, Mansouri et.al [61] and Mansouri & Nasseri [70].
@@
In this work, we investigate the dynamics of inflation in a fractal cosmological background  [^3] Although our framework introduces a fractional length scale $L$ and modifies the effective momentum structure (see Sec.4), this work should be clearly distinguished from fractional cosmology [49, 51, 95, 57, 65, 33, 98, 59, 97, 73, 23, 39, 44, 87] based on fractional calculus or non-local fractional operators [81, 91, 50]. In particular, we do not employ non-local fractional derivatives (as compared to say works which use Riesz’s fractional derivative [57, 13, 80]) in the action [41, 43, 59] and model cosmology with a Orstein-Uhlenbeck like fractional differential equation [42]. Our approach is purely phenomenological and geometric, encoding derivatives from standard FLRW cosmology through an effective dimension $D$ and horizon thermodynamics. . By deriving modified slow-roll parameters, analyzing their evolution under different potentials (Cubic, Starobinsky ($R+R^{2}$) [99] and Natural [48, 47, 35]), and extending the perturbation theory to include fractal modifications, we aim to assess the viability of fractal inflationary models. In particular, we explore how the fractal dimension influences the duration of inflation, the behavior of perturbations, and the resulting predictions for the scalar spectral index [37, 83, 103, 27, 93, 72, 102]. Our results allow comparison with observational data, providing a means to constrain the fractal dimension and to evaluate the physical consistency of inflation within a fractal cosmology. Throughout our manuscript, we have treated the fractal dimension $D$ as a constant. This contrasts with models such as that of Abdalla, Afshordi et.al [1] in which, $D$ evolves dynamically from near homogeneity ($D\simeq 3$) to a smaller asymptotic value, yielding a self-similar structure only in the late universe.
@@
2 A Review Of Fractal Cosmology
Before studying inflation in a fractal setup, it is necessary to establish the background cosmological theory in such a geometry. In fractal cosmology [55], the large-scale structure of spacetime is described by an effective dimension $D$, introducing corrections to the standard definitions of horizon area, entropy, and expansion dynamics [55]. In what follows, we review these modifications and derive the corresponding forms of the Friedmann and continuity equations that will form the basis of our inflationary analysis.
In standard cosmology, spacetime is treated as a smooth, differentiable manifold of integer dimension [107]. However, in a fractal framework, this assumption is relaxed, and the effective dimensionality of spacetime becomes scale-dependent, reflecting the possibility that the microscopic structure of the universe may not be strictly three-dimensional. Various approaches to quantum gravity and spacetime discreteness, such as causal dynamical triangulations [5, 76] and asymptotic safety [64, 17, 38], suggest a dynamical reduction of the spectral dimension at small scales. Fractal cosmology provides a phenomenological framework for capturing such effects by modifying geometric quantities such as area, volume, and horizon entropy via an effective dimension $D\not=3$. These corrections can alter the Friedmann equations and, in turn, the inflationary dynamics.
To begin with, recall that the line element of the FLRW universe [107] is given by
@@ 2 A Review Of Fractal Cosmology
There are several nonequivalent ways to implement a fractal modifications of cosmology. The work of G. Calcagni [21] proposed that phenomenological fractal measures and Lebesgue-Stieltjes integration which leads to a modified scaling of area/volume. There have been works where fractal-like spectral-dimension flows emerge in discrete quantum gravity approaches such as causal dynamical triangulations and asymptotic safety [64, 17, 38, 5, 76]. Furthermore, there have been thermodynamic derivations where the horizon entropy/area relations are deformed and the Friedmann equations follow from the first law on the apparent horizon (see for eg. [55]). In this work, we adopt the thermodynamic/cosmological horizon route as our working model.
@@
where $L$ denotes a characteristic fractional length scale associated with the underlying fractal microstructure. Throughout this work, we treat $L$ as a fixed microscopic scale of order Planck length. The factor $(4\pi)^{(D-3)/3}$ ensures that the standard FLRW relations are recovered in the limit $D\rightarrow 3$.
@@ 3 Fractal Inflation Dynamics
Here $V(\phi)$ is the inflaton potential, and this inequality generalizes the standard slow-roll condition to a fractal spacetime. Eqn. (19) therefore constrains the allowed values of $D$ such that,
$D>\sqrt{3}$ (20)
ensuring that $\dot{\phi}$ remains real.
@@
Substituting eqns.(16) and (17) into the continuity eqn.(14), we obtain the modified Klein-Gordon equation for the inflaton field:
@@
The presence of $D$ in the friction term indicates that the rate of damping of the inflaton’s motion depends directly on the effective fractal dimension; a decrease in $D$ reduces the Hubble friction, allowing the inflaton field to roll down the potential more rapidly and thereby shortening the duration of inflation.
@@
$\ddot{H}=-\frac{4\pi GD^{2}}{9}\left(\frac{2GL^{2}\rho}{3}\right)^{\frac{D}{3}-1}\left[\begin{split}\left(\frac{D}{3}-1\right)\frac{(-DH\dot{\phi}^{2})}{\frac{\dot{\phi}^{2}}{2}+V(\phi)}\\ -\dot{\phi}(DH\dot{\phi}+V_{,\phi})\end{split}\right].$ (24)
@@
A sustained slow-roll inflation requires both $|\epsilon_{1}|<<1$ and $|\epsilon_{2}|<<1$ (note that $|\epsilon_{2}|<<1$ is a condition for slow-roll inflation, not inflation in general). The fractal background of the universe modifies $\epsilon_{2}$ to,
$\begin{split}\epsilon_{2}=\frac{D\dot{\phi}^{2}}{3\left[\frac{\dot{\phi}^{2}}{2}+V(\phi)\right]}[4-D]-2D-\frac{2V_{,\phi}}{H\dot{\phi}}=2\epsilon_{1}[4-D]-2D-\frac{2V_{,\phi}}{H\dot{\phi}}\\ =2\epsilon_{1}[4-D]-2D+\frac{2(DH\dot{\phi}+\ddot{\phi})}{H\dot{\phi}}=2\epsilon_{1}[4-D]+\frac{2\ddot{\phi}}{H\dot{\phi}}.\end{split}$ (26)
Note that in the Standard $D=3$ case, the above expression reduces to the well known form
$\epsilon_{2}=2\epsilon_{1}-6+\frac{2(3H\dot{\phi}+\ddot{\phi})}{H\dot{\phi}}=2\epsilon_{1}+\frac{2\ddot{\phi}}{H\dot{\phi}}.$ (27)
@@
(a) Evolution of $\epsilon_{1}$ with the number of e-folds.
Refer to caption Refer to caption
(b) Evolution of $\epsilon_{2}$ with a zoomed region near $N=0$.
Figure 1: Evolution of the slow-roll parameters for the linear potential $V=V_{0}\left(1+\frac{\phi}{10}\right)$ for four values of the fractal dimension $D=2,3,4$. The initial conditions are $\phi=10^{-2}$ and $\dot{\phi}=0$. Panel (a) shows the first slow-roll parameter $\epsilon_{1}$, while panel (b) shows $\epsilon_{2}$ with an inset highlighting the early evolution for $0\leq N\leq 0.01$.
(a) Evolution of $\epsilon_{1}$ with the number of e-folds.
Refer to caption Refer to caption
(b) Evolution of $\epsilon_{2}$ with a zoomed region near $N=0$.
Figure 2: Evolution of the slow-roll parameters for the cubic potential $V=V_{0}\left(1+\frac{\phi^{3}}{10}\right)$ for four values of the fractal dimension $D=2,3,4$. The initial conditions are $\phi=10^{-2}$ and $\dot{\phi}=0$. Panel (a) shows the first slow-roll parameter $\epsilon_{1}$, while panel (b) shows $\epsilon_{2}$ with an inset highlighting the early evolution for $0\leq N\leq 0.01$.
[…]
Having analyzed monomial potentials such as cubic and linear, we now turn our attention to the Starobinsky ($R+R^{2})$ model of inflation [99], which arises from a purely geometric modification of gravity through the inclusion of an $R^{2}$ term in the action. Since this model is both theoretically well-motivated and tightly constrained by current CMB data, it provides an ideal benchmark for testing the robustness of the fractal framework against high-precision cosmological observations.
@@
$S=\frac{M_{Pl}^{2}}{2}\int d^{D+1}x\sqrt{-g}\left(R+\frac{R^{2}}{6M^{2}}\right).$ (30)
@@
$V(\phi)=\frac{3}{4}M^{2}M_{Pl}^{2}\left(1-e^{-\sqrt{\frac{D-1}{D}}\frac{\phi}{M_{Pl}}}\right)^{2}.$ (31)
@@
$\epsilon_{1}\simeq\frac{2(D-1)}{D}\frac{x^{2}}{(1-x)^{2}}\left(\frac{2GL^{2}\rho}{3}\right)^{1-\frac{D}{3}}.$ (32)
@@
$N\simeq\frac{D}{2(D-1)x}\left(\frac{2GL^{2}\rho}{3}\right)^{\frac{D}{3}-1}.$ (33)
@@
$\epsilon_{1}=\frac{D}{2(D-1)N^{2}}\left(\frac{2GL^{2}\rho}{3}\right)^{1-\frac{D}{3}},$ (34)
@@
$\begin{split}\epsilon_{2}=\frac{4x}{(1-x)^{2}}\frac{D-1}{D}\left(\frac{2GL^{2}\rho}{3}\right)^{1-\frac{D}{3}}+\left(\frac{D}{3}-1\right)\frac{DH\dot{\phi}^{2}}{V(\phi)}\\ =\frac{2}{N}+\left(\frac{D}{3}-1\right)\frac{DH\dot{\phi}^{2}}{V(\phi)}.\end{split}$ (35)
@@
In the fractal framework, the slow-roll parameter $\epsilon_{1}$ receives an additional multiplicative suppression $\frac{D-1}{D}\left(\frac{2GL^{2}\rho}{3}\right)^{\frac{D}{3}-1}$ controlled by the effective dimension $D$, as shown in eqn.(34). For $\frac{2GL^{2}\rho}{3}<1$ (which is quite realistic for microscopic values of $L$) and $D<3$, this gravitational correction reduces $\epsilon_{1}$ and $|\epsilon_{2}|$ independently of the flattening of the inflationary potential. As a result, the unique phenomenological advantage of plateau models such as Starobinsky inflation becomes less pronounced: slow roll can be achieved without relying exclusively on extreme potential flatness. This leads to a partial degeneracy between potential-driven and geometry driven suppression mechanisms, rendering plateau models less uniquely preferred, though still fully consistent within the fractal cosmology framework. This particular result goes in hand with the recent CMB and BAO data [8], providing a fractal perspective to the matter.
Note that, since the Higgs inflationary model [26] is also a plateau model like Starobinsky, we similarly expect that Higgs inflation is also less preferred in a fractal universe, once again going in hand with the recent CMB and BAO data [8]. A full analysis regarding the RG flow of $\lambda$ [96] and the subsequent reheating [94] will be carried forward in future works.
We now turn to a qualitatively different class of inflationary scenarios, namely Natural Inflation [48, 47, 35]. In contrast to the Starobinsky potential, which originates from higher curvature corrections to gravity, Natural Inflation is driven by a pseudo Nambu Goldstone boson [105] whose flat potential is protected by an underlying shift symmetry.
@@
Let $f$ be the axion decay constant, which serves to measure the extent of symmetry breaking and let $\Lambda$ set the energy scale for the symmetry breaking. The scalar potential for Natural Inflation is then given by a periodic form arising from the spontaneous breaking of a global symmetry, and is expressed as
@@
Hence, in the natural inflationary case, the only modification sourced by the fractal dimension $D$ is that of factor $\left(\frac{2GL^{2}\rho}{3}\right)^{1-\frac{D}{3}}$.